What is the density of a substance that takes up 3.4e3 cubic cm and weighs 1.96 kg?
Answer #1

Density = Mass/ Volume

Mass = 1.96 kg = 1.96 kg * (1000 g/ 1 kg) = 1.96 * 103 g

Volume = 3.4 * 103 cm3

So, Density = 1.96 * 103 g/ 3.4 * 103 cm3

Or, Density = 0.58 g/ cm3